# COPAL: Continual Pruning in Large Language Generative Models

### v1

@@ 6 Conclusion
In this paper, we introduced the concept of continual pruning that addresses pruning under a continual model adaptation setting, bypassing the requirement for model re-training. Two common problems observed in continual pruning, weight stasis and forgetting, were addressed through our proposed COPAL framework. We presented mathematical background, formulations, and full derivations of COPAL. Throughout the comprehensive evaluation, COPAL significantly outperformed existing methods in both Backward Transfer (BWT) reduction and Perplexity (PPL) performance, showcasing remarkable consistency and adaptability across various pruning structures, datasets and models, including LLAMA-7B, LLAMA-30B, and LLAMA-65B.
7 Broader Impact
@@ C.2 Moore-Penrose Pseudoinverse
Appendix D Sum of Absolute Gradients
By taking the absolute value of each term in the summation, we transform the equation to reflect this emphasis on magnitude:
$\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}=\sum_{i=0:k}\sum_{j}2\left|d\mathbf{y}^{i}_{j}\frac{\partial f}{\partial\mathbf{W}}\right|$ (28)
For simplicity, we separate the terms for the current dataset $k$ from the previous datasets (up to $k-1$):
$\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}=\sum_{j}2\left|d\mathbf{y}^{k}_{j}\frac{\partial f}{\partial\mathbf{W}}\right|+\sum_{i=0:k-1}\sum_{j}2\left|d\mathbf{y}^{i}_{j}\frac{\partial f}{\partial\mathbf{W}}\right|$ (29)
Recognizing that the second term in Equation (4) represents the sum of absolute values of gradients for all datasets up to $k-1$, we can rewrite it as $\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}_{k-1}$. Therefore, we arrive at:
$\begin{split}\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}&=\sum_{j}2\left|d\mathbf{y}^{k}_{j}\right|\left|\frac{\partial f}{\partial\mathbf{W}}\right|+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}_{k-1}\\ &=\nabla^{\prime}_{d\mathbf{W}}\tilde{\mathcal{L}}^{k}+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k-1}\end{split}$ (30)
where ~$\tilde{\mathcal{L}}^{k}$ is loss function for only dataset k.
Appendix E Directional Derivative
The directional derivative $D_{\mathbf{v}}f$ of a function $f(\mathbf{x})$ along a direction $\mathbf{v}$ is defined as:
$D_{\mathbf{v}}f(\mathbf{x})=\nabla f(\mathbf{x})\cdot\mathbf{v}$ (31)
where $\nabla f(\mathbf{x})$ is the gradient of $f$ and $\cdot$ is the dot product. It measures how much $f$ will change if we move from $\mathbf{x}$ in the direction $\mathbf{v}$.
We use magnitude of directional derivative of $\mathcal{L}^{i}_{j}$ along $\mathbf{W}$ as our measure of importance weights denoted by $\mathbf{W}_{k}^{*}$. ~$\tilde{\mathcal{L}}^{k}$ is loss function for only dataset $k$.
$\begin{split}\mathbf{W}^{*}_{k}&=\sum_{i=0:k}\sum_{j}\left|D_{\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\sum_{i=0:k}\sum_{j}\left|\mathbf{W}\cdot\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\left|\mathbf{W}\right|\cdot\sum_{i=0:k}\sum_{j}\left|\nabla_{d\mathbf{W}}\mathcal{L}^{i}_{j}\right|\\ &=\left|\mathbf{W}\right|\cdot\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}^{k}\\ &=\left|\mathbf{W}\cdot\right|(\nabla_{d\mathbf{W}}\tilde{\mathcal{L}}^{k}+\nabla^{\prime}_{d\mathbf{W}}\mathcal{L}_{k-1})\\ &=\sum_{j}\left|\mathbf{W}\cdot\nabla_{d\mathbf{W}}{\mathcal{L}}^{k}_{j}\right|+\mathbf{W}^{*}_{k-1}\end{split}$ (32)
Appendix F Automating threshold for rsparsity
@@
Appendix G Weight Stasis inequality detailed proof
In the context of calibration-guided pruning strategies, a rigorous mathematical justification is necessary to explain the transition of the inequality to the binary mask $\mathcal{M}_{i}$. We present a formal argument to elucidate this phenomenon.
We consider the importance of weights, $\mathbf{W}^{*}_{i}=\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i}\right|$ as the basis for the pruning process, where $\mathcal{R}_{i}$ is scaling or ranking the weights $\mathbf{W}_{i}$ with some criteria. For a given dataset $i$, mask ${M}_{i}$, and $\mathcal{I}$ as indicator function, the pruned weight matrix $\mathbf{W}^{p}_{i}$ is obtained as follows:
Upon transitioning from dataset $i$ to dataset $i+1$, the pruning mechanism starts with $\mathbf{W}_{i+1}=\mathbf{W}^{p}_{i}$ as the initial weight matrix. This initial matrix already has a set of pruned (zeroed-out) weights. These zeroed weights in $\mathbf{W}_{\text{i+1}}$ will not have any effect for the new dataset $i+1$ inputs due to their suppressed influence, shown below:
$\begin{split}\mathcal{M}_{i+1}&=\mathcal{I}\left(\mathbf{W}^{*}_{i+1}<\mathcal{T}_{s}\right)\\ &=\mathcal{I}(\left|\mathbf{W}_{i+1}\cdot\mathcal{R}_{i+1}\right|<\mathcal{T}_{s})\\ &=\mathcal{I}(\left|(\mathbf{W}_{i}\cdot\mathcal{M}_{i})\cdot\mathcal{R}_{i+1}\right|<\mathcal{T}_{s})\\ &=\mathcal{I}(\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i+1}\right|\cdot\left|\mathcal{M}_{i}\right|<\mathcal{T}_{s}))=\begin{cases}0&\text{if }m<\mathcal{T}_{s},m\in\left|\mathbf{W}_{i}\cdot\mathcal{R}_{i+1}\right|\cdot\left|\mathcal{M}_{i}\right|\\ 1&\text{otherwise.}\end{cases},\\ &=\mathcal{M}_{i}\\ \mathbf{W}_{i+2}&=\mathbf{W}_{i+1}=\mathbf{W}_{i}\cdot\mathcal{M}_{i}\end{split}$ (35)
Appendix H Overall standard deviations
Table 4: Results of overall mean and standard deviation for continual pruning on wikitext2, ptb, c4 datasets with all incremental permutations on LLAMA-7B,30B,65B. Unstructured pruning with $50\%$ sparsity ratio.
@@ Appendix H Overall standard deviations
Appendix I Baseline Details
Magnitude pruning, a method detailed by Zhu & Gupta (2017), effectively compresses models by removing weights with the smallest absolute values, which are deemed least impactful on the network’s output. In contrast, SparseGPT, introduced by Frantar & Alistarh (2023), integrates sparsity into the post-training process of transformer-based models. It often employs methods such as utilizing the Hessian matrix to identify weights that can be pruned with minimal loss in performance, effectively making the model sparse from the outset. Finally, WANDA, as proposed by Sun et al. (2023), does pruning by analyzing weight and activation distributions. This technique identifies and prunes network segments that minimally contribute to outputs.

### v2

@@ 6 Conclusion
In this paper, we introduced the concept of continual pruning that addresses pruning under a continual model adaptation setting, bypassing the requirement for model re-training. Two common problems observed in continual pruning, weight stasis and forgetting, were addressed through our proposed COPAL framework. We presented mathematical background, formulations, and full derivations of COPAL. Throughout the comprehensive evaluation, COPAL significantly outperformed existing methods in both Backward Transfer (BWT) reduction and Perplexity (PPL) performance, showcasing remarkable consistency and adaptability across various pruning structures, datasets and models, including LLaMA-7B, 13B, 30B, 65B.
Impact Statement
@@ C.2 Moore-Penrose Pseudoinverse
Appendix D Automating threshold for rsparsity
@@
Appendix E Overall standard deviations
Table 4: Results of overall mean and standard deviation for continual pruning on wikitext2, ptb, c4 datasets with all incremental permutations on LLaMA-7B, 13B, 30B, 65B. Unstructured pruning with $50\%$ sparsity ratio.
@@ Appendix E Overall standard deviations
LLAMA-13B
Dense (no pruning) - 6.990±1.634
Magnitude WS 28.935 ± 9.026
SparseGPT 0.606 ± 0.381 8.467 ± 1.766
Wanda 0.203 ± 0.065 8.570 ± 1.893
COPAL (ours) 0.029 ± 0.028 8.354 ± 1.851